### v1

# IEEEexample:BSTcontrol

Abstract
The advent of internet medicine provides patients with unprecedented convenience in searching and communicating with doctors relevant to their diseases and desired treatments online. However, the current doctor recommendation systems fail to fully ensure the professionalism and interpretability of the recommended results. In this work, we formulate doctor recommendation as a ranking task and develop a large language model (LLM)-based pointwise ranking framework. Our framework ranks doctors according to their relevance regarding specific diseases-treatment pairs in a zero-shot setting. The advantage of our framework lies in its ability to generate precise and explainable doctor ranking results. Additionally, we construct DrRank, a new expertise-driven doctor ranking dataset comprising over 38 disease-treatment pairs. Experiment results on the DrRank dataset demonstrate that our framework significantly outperforms the strongest cross-encoder baseline, achieving a notable gain of +5.45 in the NDCG@10 score while maintaining affordable latency consumption. Furthermore, we comprehensively present the fairness analysis results of our framework from three perspectives of different diseases, patient gender, and geographical regions. Meanwhile, the interpretability of our framework is rigorously verified by three human experts, providing further evidence of the reliability of our proposed framework for doctor recommendation.
{IEEEkeywords}
Doctor recommendation, explainable recommendation, large language model, text ranking.
1 Introduction
[…]
The rapid development of internet medicine provide patients with convenient access to trans-regional disease diagnosis and treatment services [1]. Through professional online healthcare platforms, patients can search for and select experienced doctors based on their potential diseases and required treatments, including medication, surgery, and radiotherapy. Although some authoritative leaderboards have been released, such as the China Hospital and Departments Reputation Ranking [^1] https://rank.cn-healthcare.com/fudan/national-general, this leaderboard can only guide patients in choosing from a limited number of hospitals and departments, let alone specific doctors. For patients lacking medical knowledge, accurately selecting a doctor who matches their condition remains challenging. To provide high-quality online medical services, a crucial task is to recommend an appropriate doctor from a large pool of candidates for a specific patient [2, 3].
Recent studies have shown significant interest in developing doctor recommendation systems. Early methods rely on basic statistical indicators such as the number of patient reviews and treatment cases [4, 5]. These methods fail to fully capture the nuanced professional expertise of doctors, consequently compromising the accuracy of rankings. Current methods seek to improve doctor recommendation through machine learning techniques, but some significant limitations remain. Machine learning methods usually require sufficient labeled data for model training, which poses a significant challenge in data-sparse medical scenarios due to the high cost and effort of data annotation. Some studies define the ground truth label in doctor-patient interactions based on whether the consultation order is successfully completed by both parties [6], or whether the doctor later responded to the patient’s inquiry [7]. In such cases, the optimization objectives of recommendation systems tend to be skewed towards commercial considerations. This commercial bias, to some extent, undermines the focus on doctors’ expertise and may not adequately address patients’ needs to find experts in specific medical fields. Additionally, explanations are crucial in ensuring the interpretability of recommendation systems [8], particularly in the medical context of doctor recommendation. However, the interpretability of doctor recommendation is still largely unexplored at present.
Figure 1: An overview of our proposed zero-shot LLM-based doctor ranking framework. We leverage a pointwise ranking method to efficiently and effectively evaluate the professional relevance of candidate doctors in addressing specific patient medical needs. By utilizing fine-grained relevance labels and incorporating query-specific ranking criteria, our framework ensures consistent and accurate doctor evaluations. Furthermore, it provides clear and structured doctor evaluation rationales as recommendation explanations, enhancing the interpretability and transparency of the doctor recommendation process. It should be noted that the ranking criteria and recommendation explanation shown in the figure are simplified versions.
In this work, we focus on constructing an accurate, fair, and explainable doctor recommendation system based on the professional relevance of doctors in addressing patient medical needs. We formulate doctor recommendation as a ranking task and introduce DrRank, a novel expertise-driven doctor ranking dataset that spans various disease-treatment pairs. To create comprehensive doctor profiles, we integrate multidimensional professional information, such as their introductions, areas of expertise, community involvement, awards received, and research achievements. Furthermore, we involve medical experts to annotate the professional relevance of doctors in relation to specific disease-treatment pairs. In the field of information retrieval, cross-encoder models rely on annotated relevance data for supervised training [9]. These models have exhibited impressive performance in various out-of-domain ranking tasks [10]. However, their efficacy in the specific context of doctor ranking has yet to be explored, primarily because such a task is typically absent from previous training datasets.
Large language models (LLMs), backed by extensive pre-training data, demonstrate the ability to perform text ranking tasks based on user prompts [11, 12, 13, 14]. The remarkable generalization capabilities and inherent world knowledge of LLMs suggest significant potential for the doctor ranking task. Inspired by this, we propose an LLM-based doctor ranking framework as illustrated in Fig. 1, which ranks doctors according to their relevance regarding specific disease-treatment pairs. It operates in a zero-shot setting, thus eliminating the need for vast amounts of labeled data for training Specifically, we employ a pointwise ranking method to ensure the efficiency of the doctor ranking process. We utilize fine-grained relevance labels to assist the LLM in more accurately distinguishing the nuanced differences in doctors’ relevance across various disease-treatment pairs. To enhance the consistency and accuracy of the LLM’s ranking results, we prompt the LLM to generate specific ranking criteria tailored to a specific disease and treatment option. The generated ranking criteria serve as evaluation guidelines, providing clear and transparent directives for the LLM during the ranking decision-making process. To improve the interpretability of the doctor recommendation process, once the LLM has generated doctor relevance predictions, we continuously prompt it to further generate evaluation rationales as the explanations for doctor recommendation. We observe that the stochastic behavior of the LLM can potentially influence both the relevance generation and rational generation stages. To address this, we propose a forced elicitation mechanism that ensures the LLM consistently generates the desired results.
Experimental results demonstrate that LLMs, with a scale exceeding 7 billion parameters, exhibit superior performance compared to cross-encoder models for doctor ranking task, particularly when ranking criteria are introduced. Additionally, we present the time costs associated with utilizing the LLM for pointwise prediction and the generation of evaluation rationales, illustrating the feasibility of our framework in real-world applications. Furthermore, we analyze the distribution characteristics of the LLM prediction results and ground truth relevance labels, revealing how our framework works in practice. Meanwhile, we evaluate the fairness and interpretability of our framework. Experimental results show that our framework exhibits better fairness performance compared to cross-encoder models across a range of disease-treatment pairs. However, we also acknowledge concerns regarding potential gender biases. Through manual evaluation of the ranking criteria and evaluation rationales within our framework, we highlight the significant potential and practical utility of our framework in serving patients.
The main contributions of our work can be summarized as follows:
  1. 1.
    We construct a new expertise-driven dataset for objectively evaluating models in doctor ranking.
  2. 2.
    We develop a zero-shot LLM-based doctor ranking framework for doctor recommendation, addressing the issue of limited labeled data for model training.
  3. 3.
    Through extensive experiments, we demonstrated the accuracy, fairness, and interpretability of our framework, and we have released our code to ensure reproducibility.
2 Related Work
2.1 Doctor Recommendation
Doctor recommendation aims to help patients find the most suitable doctors. While many works in recommendation focus on modeling target users from historical behavior [15], doctor recommendation can only rely on limited information in patients’ chief complaints for modeling due to privacy concerns. Modeling doctors effectively is another pivotal area of research. Some studies focus on constructing doctor profiles based on basic features available on online healthcare platforms [16, 17], such as years of experience, pricing, and online reviews. Furthermore, dialogue data between doctors and patients has been utilized to enhance doctor modeling and to investigate the combined effects of doctor profiles and previous dialogues with other patients [6, 7]. Additional research efforts are directed towards developing personalized recommendation to help patients find doctors who align with their personal preferences [18, 19]. Che et al. [20] delve into the challenge of integrating data of varying structures across multiple medical platforms, exploring innovative ways to leverage this diverse information effectively. In this work, we focus on doctor recommendation based on comprehensive doctor profiles and leverage the zero-shot ranking capabilities of LLMs to escape the predicament of training a doctor ranking model with scarce labeled data.
2.2 Large Language Models in Zero-shot Text Ranking
Shifting from tuning-based learning to rank driven by cross-encoder models [9, 10] such as MonoT5 [21] and RankT5 [22], there is an emerging line of research exploring how to utilize general-purpose LLMs for zero-shot text ranking in out-of-domain scenarios. The current LLM-based ranking methods can be mainly divided into the following three categories: (1) Pointwise strategy utilizes LLMs to measure the relevance between a query and a single passage by query generation [23] or relevance generation [24, 11]. (2) Pairwise strategy prompts LLMs to determine which of two passages is more relevant to a query and aggregates a final relevance score for each passage from all pair comparisons, where efficient sorting algorithms, such as heapsort, can be used to speed up the ranking process particularly in top-k ranking [13]. Setwise strategy enhances the efficiency of pairwise comparisons by selecting the most relevant passage from multiple candidates [14]. (3) Listwise strategy instructs LLMs to generate a permutation of passages, arranged in descending order of relevance to a query [12]. Due to the limited input length of LLMs, a sliding window approach is used to rerank a subset of candidate passages each time, moving the window from back to front to cover all candidates. In this work, we formulate doctor recommendation as a ranking task and develop a zero-shot LLM-based pointwise ranking framework which focuses on the accuracy, consistency and interpretability of doctor ranking.
3 DrRank Dataset
In this section, we elaborate on the construction of the DrRank dataset and provide an overview of its statistical information.
Figure 2: An example of doctor profiles used in the DrRank dataset, highlighting the nine key fields. This detailed profile format ensures comprehensive representation of a doctor’s professional background and expertise.
3.1 Data Collection and Preprocessing
The doctor profiles used in this study is primarily sourced from Haodf Online [^2] https://www.haodf.com, one of the biggest online healthcare platforms in China. Each doctor profile comprises 9 fields: “Title”, “Specialty”, “Affiliation”, “Department”, “Introduction”, “Expertise”, “Social Service”, “Awards”, and “Research”, as depicted in Fig. 2. Over 600,000 doctor profiles were collected in total, forming a comprehensive resource for analysis. To proceed, we first selected a range of common cancers and critical illnesses, along with their typical treatment options, resulting a set of 38 disease-treatment pairs. For each pair, such as lung cancer and surgical treatment, we employed the open-sourced Alibaba-NLP/gte-multilingual-reranker-base [^3] https://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base model [25] to assess the relevance of the introductions from the doctor profiles in our extensive pool to the disease. This process yielded an initial list of 100 candidate doctors for each disease-treatment pair.
Figure 3: Cumulative distribution functions (CDF) of doctor profile lengths from two datasets: DrRank and Collected Doctor Pool. The x-axis represents the length of tokenized doctor profiles, and the y-axis represents the cumulative percentage of sequences. The vertical dashed lines at 1024, 2048, and 4096 tokens indicate key truncation thresholds. Each of these thresholds is annotated with the percentage of the DrRank dataset that is covered up to that token count.
3.2 Data Annotation
For each disease-treatment pair, two experts in this field were interviewed for approximately one hour. They evaluated the relevance of the doctors on our initial list and also supplemented it with additional professional doctors who were missing from the list but should be included based on the actual situation. Each doctor on the list was assigned a professional relevance score ranging from 0 to 5, with 5 indicating the highest relevance and 0 representing irrelevance. Doctors who received consistent relevance scores from both experts were retained directly. In cases of disagreement between the two experts regarding the assigned relevance score, a third independent expert was consulted to provide the final evaluation.
3.3 Hard Negative Mining
We considered doctors with a relevance score greater than 1 as positive samples and those with a relevance score of 0 as negative samples. To balance the number of positive and negative samples for each disease-treatment pair and enhance the complexity of distinguishing between positive and negative samples, we conducted hard negative mining. Through statistical analysis of the doctor profiles of annotated positive samples, we found that they tend to be relatively long. Therefore, we filtered the doctor profiles to include only those with introductions exceeding a minimum length of 1,024  [^4] Unless otherwise specified, all lengths mentioned in this work are defined based on the tokenizer inherent to the Qwen2.5 series.. This filtering process resulted in a subset of 5,000 unique doctors with sufficiently rich profile information. For each disease-treatment pair, we utilized the gte-multilingual-reranker-base [25] model once again to rank the filtered set of doctors based on the relevance score between the disease and the doctor’s introduction. To minimize false negatives, we excluded the top 1% of doctors from the ranked list. Then, from the remaining list, we iteratively added negative samples not initially included, in descending order of relevance, until we achieved a 1:1 ratio of positive to negative samples for each disease-treatment pair. To further introduce more challenging negative samples, we randomly replaced 30% of the previously selected negative samples with positive samples from other pairs. Specifically, we regarded the positive samples with relevance scores of 4 and 5 from other disease-treatment pairs as negative samples for the current pair. Finally, we conducted a manual review to ensure that the all selected negative samples for each disease-treatment pair were indeed unrelated to it.
Figure 4: Heatmap of the annotated professional relevance distribution across positive samples for different disease-treatment pairs in the DrRank dataset. The x-axis represents the professional relevance (ranging from 1 to 5), while the y-axis lists various disease-treatment pairs.
3.4 Dataset Statistics
The DrRank dataset, constructed through the aforementioned process, includes 30 disease types, comprising 23 cancers and 7 other critical illnesses. Each disease is linked to one or more treatment options such as medication, surgery, radiotherapy, and, in the case of leukemia, Hematopoietic Stem Cell Transplantation (HSCT). Although most diseases are associated with a single treatment option, some involve multiple treatments, resulting in a total of 38 unique disease-treatment pairs. Our dataset includes 4325 doctor profiles, of which 2,593 are unique. Some doctors are listed in multiple disease-treatment pairs due to their expertise spanning a variety of conditions and treatments. Fig. 4 illustrates the unbalanced distribution of annotated professional relevance scores among positive samples across all the disease-treatment pairs in the DrRank dataset. Furthermore, Fig. 3 reveals substantial variability in the length of doctor profiles, with the average length in the DrRank dataset exceeding that of our initial doctor pool. These findings underscore the complexity of the doctor ranking task in real-world scenarios, given the significant variation in the volume and diversity of available information. Looking ahead, we plan to expand the dataset by incorporating more disease-treatment pairs.
4 LLM-based Doctor Ranking Framework
4.1 Problem Formulation
[…]
Definition 1 (Doctor Recommendation)
Given a patient’s medical need $q\in\mathcal{Q}$ and a candidate doctor $d\in\mathcal{D}$, the goal of doctor recommendation is to evaluate the professional relevance of doctors to patients’ specific medical needs, i.e., $f(q,d)\rightarrow\mathbb{R}$, which serves as the basis for doctor ranking.
4.2 LLM-based Doctor Ranking
Fig. 1 presents a comprehensive overview of our proposed doctor ranking framework, which integrates several key components to ensure accurate, efficient, and explainable doctor recommendation. Below, we will elaborate on the details of our framework.
4.2.1 Exploiting Efficient and Effective Pointwise Ranking
Although pairwise and listwise ranking methods have the capability to compare candidates relatively during inference time, their practical application in real-world scenarios remains extremely challenging. Pairwise comparisons suffer from quadratic complexity [13], incurring significant computational overhead. Meanwhile, the sliding window mechanism employed by listwise method is inherently difficult to parallelize [12], resulting in substantial latency in large-scale medical data sorting tasks. Additionally, both pairwise and listwise methods are susceptible to position bias [26, 27, 13], where the order of candidates can influence the ranking outcomes. This bias introduces instability, severely compromising the credibility of ranking results, particularly in critical medical scenarios where precision and reliability are paramount. In contrast, the pointwise ranking method excels with its high degree of parallelism, allowing for independent scoring of each candidate [23, 24]. A notable advantage of the pointwise method lies in the fact that it generates scores which are free from position bias. This ensures that the scores are stable, facilitating easy reproducibility of ranking results. Furthermore, by harnessing the robust language generation capabilities of LLMs, generating the scoring rationale for each candidate becomes straightforward, enhancing the transparency and trustworthiness of the ranking process. In comparison, pairwise and listwise methods usually focus on relative comparisons between pairs or lists of candidates, making it difficult to extract and present individual, easily understandable rationales for each candidate. Considering the aspects of parallel efficiency, score stability and interpretability mentioned above, the pointwise ranking method emerges as the ideal choice for doctor ranking.
4.2.2 Utilizing Fine-grained Labels to Optimize Distinction Relevance
Existing research utilizing LLMs as zero-shot pointwise rankers can be broadly categorized into two approaches: relevance generation [24, 11] and query generation [23]. This work focuses on the relevance generation approach. In this approach, the LLM is prompted to determine whether a document is relevant to a query, using the logits of “Yes” and “No” tokens to derive a relevance score. However, studies involving human subjects have demonstrated that binary options can sometimes lead to biased responses, whereas offering options with a higher level of granularity can produce more dependable outcomes [28]. Zhuang et al. [11] suggest incorporating detailed relevance labels into the prompt for pointwise LLM rankers, allowing them to better distinguish among documents with differing degrees of relevance to the query and thus generate a more precise ranking. Inspired by this, we employ a set of 5-level graded labels [“Top”, “High”, “Mid”, “Low”, “Not Relevant”] to assess the professional relevance of doctors in addressing patients’ specific medical needs related to a particular disease and its treatment options. For each query-doctor pair $(q,{d_{i}})$, we instruct the LLM to evaluate their professional relevance by selecting from the provided graded labels. We can obtain the corresponding $logit_{k}$ for each graded label $l_{k}$ from the LLM output as (1), which represents the raw confidence values of evaluation.
$logit_{i,k}=\mathrm{LLM}(l_{k}|q,d_{i},prompt),\;k\in[0,1,2,3,4]$ (1)
We can use these logits to derive the ranking score by calculating the expected value. Firstly, we need to assign a score $s_{k}$ to each graded label $l_{k}$ Through empirical investigation, we find that naively assigning $s_{k}=k$ (with 0 to 4 ordered from least to most relevance) already yields excellent performance. Then, we can calculate the expected scoring value as follows:
$f(q,d_{i})=\sum_{k}prob_{i,k}\cdot s_{k},$ (2)
where $\text{where }prob_{i,k}=\frac{\exp(logit_{i,k})}{\sum_{k^{\prime}}\exp(logit_{i,k^{\prime}})}.$
In this context, the marginal probability $prob_{i,k}$ can be interpreted as the likelihood of the doctor $d_{i}$’s relevance belonging to the graded label $l_{k}$. In fact, this approach to finding the expected average is akin to sampling the outputs of LLMs multiple times [11]. In preliminary experiments, we surprisingly observed that the first token output by instruction-tuned LLMs does not always match our predefined 5-level graded labels, complicating the extraction of logits for deriving ranking scores. Therefore, we propose a forced elicitation mechanism, which involves appending a predefined labeling prefix to elicit the desired graded labels from the LLM. The labeling prefix (i.e., “The professional relevance of the candidate doctor is ”) ensures that the first token output by the LLM matches our expectations perfectly.
4.2.3 Generating Ranking Criteria as Standard Guidelines
LLMs are widely known as extremely intricate “black-box” systems [29]. The iterative nature of next-token prediction results in ambiguous assessment criteria for zero-shot pointwise rankers [30], leading to a lack of consistency and transparency in the scoring process. Establishing clear criteria by which LLMs score doctors is crucial for building an accurate and explainable doctor recommendation system. Recent studies on human annotation practices have emphasized the importance of a standardized annotation guideline for ensuring clear and consistent communication with annotators [31, 32]. Inspired by this, we argue that standard ranking criteria should be provided for LLMs to ensure doctors are evaluated under a unified and transparent standard, thereby enhancing the consistency and accuracy of the final doctor ranking results. To this end, we meticulously design a prompt (as shown in Fig. 5) to instruct LLMs to generate relevant doctor ranking criteria for specific disease-treatment pairs. Initially, we repeatedly prompt LLMs to generate multiple sets of ranking criteria for a specific disease-treatment pair. Subsequently, we manually inspect these criteria sets to select the best one, based on factors such as comprehensiveness, practicality, and overall format. Using the optimal criteria as an example, we adopt a one-shot prompting approach to generate consistent ranking criteria for other disease-treatment pairs. Our method dynamically integrates query-specific (i.e., disease and treatment) ranking criteria into the ranking prompt, serving as contextual guidelines for the rankers’ decisions. These query-specific factors are essential for aligning ranking results with patients’ needs and ensuring that ranking decisions are based on the most relevant professional factors. Recently some researchers use multi-perspective criteria to further enhance pointwsie rankers [30]. This approach can be viewed as a more complex variant of our method, involving multiple rounds of LLM inference and result ensembling from each perspective. While it potentially offers enhanced performance, it results in increased complexity and computational cost. Our method, in contrast, offers a simpler yet effective solution. We verify the consistent positive impact of generated criteria on the doctor ranking performance of LLMs in our ablation experiments.
Figure 5: Prompt for ranking criteria generation.
4.2.4 Generating Evaluation Rationales as Recommendation Explanations
[…]
5 Experiments
In this section, we evaluate the proposed framework on the DrRank dataset under a zero-shot setting. Meanwhile, we present various discussions on doctor recommendation.
5.1 Experimental Settings
We omit the dataset description in this section since it has been introduced in Section 3. Other experimental settings will be described in the following parts.
Table 1: The performance and efficiency of each model on the DrRank dataset. The model sizes and maximum context lengths for cross-encoder models are indicated as [parameters, tokens]. The maximum context and generation length of Qwen2.5 series are 32K and 8K respectively. “BS” is an abbreviation for “batch size”. Scoring Time and Explanation Time are the average time consumption of each batch. Average results that are significantly (paired t-test, $p<0.05$ ) better than without ranking criteria are marked with *. The top three performances of the models with ranking criteria are highlighted in bold. The top three performances of the models without ranking criteria are underlined.
Model Running Info NDCG@10 PNR Scoring Time (s) Explanation Time (s)
Cross-Encoder
jinaai/jina-reranker-v2-base-multilingual [278M, 1K] 8 BS, 1 GPU 56.91 1.79 0.01 -
Alibaba-NLP/gte-multilingual-reranker-base [306M, 8K] 8 BS, 1 GPU 71.96 4.01 0.01 -
BAAI/bge-reranker-v2-m3 [568M, 8K] 8 BS, 1 GPU 69.26 3.42 0.02 -
LLM-based Pointwise Ranker
General-purpose LLM
Qwen2.5-0.5B-Instruct w/ Ranking Criteria 5 BS, 1 GPU 49.85 1.68 0.02 6.36
Qwen2.5-0.5B-Instruct w/o Ranking Criteria 10 BS, 1 GPU 56.21 1.94 0.02 5.22
Qwen2.5-1.5B-Instruct w/ Ranking Criteria 15 BS, 1 GPU 62.33 3.38 0.09 13.82
Qwen2.5-1.5B-Instruct w/o Ranking Criteria 15 BS, 1 GPU 62.04 3.28 0.11 9.49
Qwen2.5-3B-Instruct w/ Ranking Criteria 4 BS, 1 GPU 67.48 4.23 0.11 13.4
Qwen2.5-3B-Instruct w/o Ranking Criteria 5 BS, 1 GPU 66.95 4.12 0.11 8.55
Qwen2.5-7B-Instruct w/ Ranking Criteria 6 BS, 1 GPU 74.35* 5.31* 0.23 22.62
Qwen2.5-7B-Instruct w/o Ranking Criteria 8 BS, 1 GPU 72.68 4.9 0.16 10.51
Qwen2.5-7B-Instruct w/ Random Ranking Criteria 8 BS, 1 GPU 72.52 4.6 0.23 23.22
Qwen2.5-14B-Instruct w/ Ranking Criteria 8 BS, 2 GPUs 75.63* 6.87* 1.48 32.63
Qwen2.5-14B-Instruct w/o Ranking Criteria 12 BS, 2 GPUs 73.56 6.21 1.73 28.68
Qwen2.5-14B-Instruct w/ Random Ranking Criteria 8 BS, 2 GPUs 75.51 5.27 1.46 28.8
Qwen2.5-32B-Instruct w/ Ranking Criteria 8 BS, 4 GPUs 77.41 6.13 0.1 60.21
Qwen2.5-32B-Instruct w/o Ranking Criteria 8 BS, 4 GPUs 77.19 6.1 0.1 36.42
Qwen2.5-32B-Instruct w/ Random Ranking Criteria 8 BS, 4 GPUs 76.31 5.44 0.1 57.49
Qwen2.5-72B-Instruct w/ Ranking Criteria 5 BS, 3 GPUs 75.93 5.38 4.84 110.08
Qwen2.5-72B-Instruct w/o Ranking Criteria 10 BS, 4 GPUs 75.41 5.38 5.05 82.26
Medical LLM
MING-MOE-7B w/ Ranking Criteria 6 BS, 1 GPU 48.79 1.19 1.63 41.25
MING-MOE-7B w/o Ranking Criteria 10 BS, 1 GPU 45.43 1.25 2.14 34.22
5.1.1 Baseline Methods
We selected three popular open-sourced cross-encoder models as baselines, namely BAAI/bge-reranker-v2-m3555https://huggingface.co/BAAI/bge-reranker-v2-m3 [36], Alibaba-NLP/gte-multilingual-reranker-base, jinaai/jina-reranker-v2-base-multilingual [^6] https://huggingface.co/jinaai/jina-reranker-v2-base-multilingual, which have demonstrated state-of-the-art effectiveness in various out-of-domain tasks. We concatenate pairs of the patient query and detailed text constructed from the doctor profile, and then pass these concatenated pairs through the models to directly generate a relevance score. Importantly, these cross-encoder models are typically trained on general domain text ranking datasets, such as MS MARCO [37]. Since they have likely not encountered doctor ranking tasks during their training phase, we believe that the experimental results will rely entirely on the generalization ability of these models in the context of doctor ranking.
5.1.2 Implementation Details
To comprehensively evaluate the effectiveness of our LLM-based framework, we conducted experiments using the general-purpose instruction-tuned language models of the Qwen2.5 series [38], ranging from 0.5B to 72B in size. These experiments allowed us to demonstrate the relationship between model scale and zero-shot ranking performance. In addition to the Qwen2.5 series, we also evaluated MING-MOE [39], a novel Mixture-of-Expert (MOE)-based medical LLM designed to manage diverse and complex medical tasks. MING-MOE-7B is built upon the foundation of Qwen1.5-7B-Chat [^7] https://huggingface.co/Qwen/Qwen1.5-7B-Chat [40]. Our framework was implemented using PyTorch 2.5.1 and Transformers 4.44.2 based on Python 3.10.15, and executed on a Linux machine equipped with 8 × NVIDIA A800-SXM4-80GB GPUs and 256GB of memory. To ensure reproducibility of the LLM outputs, we adopted greedy sampling in all our experiments. We set a maximum length of 2,048 for the detailed text constructed from the doctor profile, truncating any content that exceeded this limit. Furthermore, we configured a maximum length of 1,024 for the model-generated criteria and a maximum length of 512 for the evaluation rational. The ranking criteria for all disease-treatment pairs were generated offline by the Qwen2.5-7B-Instruct model.
5.1.3 Evaluation Metrics
In our experiments, we employed the Normalized Discounted Cumulative Gain@10 (NDCG@10) [41] and Positive-Negative Ratio (PNR) [42] as our principal evaluation metrics. NDCG@k is a standard listwise accuracy metric that considers the fine-grained relevance and position of documents within the top k positions in the ranked list. PNR is a pairwise metric that measures the ratio of concordant pairs to discordant pairs in a ranked list of documents. The PNR is calculated as follows:
$PNR=\frac{\sum_{i,j\in[1,N]}\mathbf{I}\{y_{i}>y_{j}\}\mathbf{I}\{f(q,d_{i})>f(q,d_{j})\}}{\sum_{i,j\in[1,N]}\mathbf{I}\{y_{i}>y_{j}\}\mathbf{I}\{f(q,d_{i})<f(q,d_{j})\}}.$ (3)
In this context, $N$ denotes the length of a ranked list, $y_{i}$ represents the ground truth label of the i-th document, and the indicator function $\mathbf{I}\{a>b\}$ takes the value 1 if $a>b$ and 0 otherwise. Due to variations in the number of doctors across different disease-treatment pairs as detailed in Section 3, we introduced the concept of “evaluation group” for a more objective assessment of the aforementioned ranking metrics. An evaluation group refers to a candidate list with a fixed length of 50, obtained through sampling without replacement from the ranking results of each query. If the original list length is less than 50, it will be directly considered as an evaluation group. To derive the final evaluation results, we calculate the average across multiple evaluation groups. Unless otherwise specified, the default number of evaluation groups is 1,000.
5.2 Zero-shot performance and efficiency
5.2.1 Ranking Performance
To demonstrate the effectiveness of our LLM-based framework, we compare Qwen2.5 series with three cross-encoder baselines, and the results are shown in Table 1. From the results, we obtain several observations.
First and foremost, the gte-multilingual-reranker-base model serves as a very strong baseline. Qwen2.5 models, with a scale exceeding 7B, exhibit superior performance compared to all baselines in terms of NDCG@10 and PNR evaluation metrics. This outstanding performance underscores LLMs’ robust domain generalization capabilities, thereby clearly demonstrating the effectiveness of our framework in assessing a doctor’s professional relevance in addressing patients’ specific medical needs. As the model scale increases, its ability to accurately rank doctors enhances gradually, although we do observe diminishing returns and a slight decline at higher scales. Small-scale LLMs exhibit poorer performance, highlighting the importance of a sufficient model scale for tackling intricate tasks like doctor ranking.
Second, the ranking criteria significantly improved the doctor ranking performance of middle-scale Qwen2.5 models. It is worth noting that the 7B model, aided by ranking criteria, achieved an NDCG@10 score of 74.35, surpassing the 14B model without ranking criteria, which scored 73.56. This underscores the significant potential of ranking criteria in enhancing LLM ranking decisions. However, we observed that for the very small 0.5B model, ranking criteria actually caused a substantial performance loss, with the NDCG@10 score dropping from 56.21 to 49.85. Meanwhile, we noticed minimal improvement effects of the ranking criteria on larger models, such as the 32B and 72B models. This suggests that although ranking criteria can enrich ranking tasks and potentially raise the upper limit of LLMs’ ranking capabilities, they also increase the minimum capacity requirements for the effective execution of the ranking tasks. Additionally, we investigated the underlying mechanism of the ranking criteria: whether it is the overall format or the specific content related to the disease-treatment pair that contributes to the ranking decisions. This question resembles the inquiry into which aspects of demonstrations contribute to the in-context learning capabilities of LLMs [43]. For each disease-treatment pair, we randomly inserted an irrelevant yet format-consistent ranking criteria into the ranking prompt. We observed that the random ranking criteria negatively impacted the 7B and 32B models. However, for the 14B model, there was still a considerable performance increase compared to the case without the ranking criteria. We argue that both the content and format of the ranking criteria likely work together to guide the ranking decisions of LLMs.
Third, we explored the potential of medical LLMs in doctor ranking using MING-MOE-7B. Although our experiments show an improvement in the NDCG@10 score due to the ranking criteria, increasing from 45.43 to 48.79, the performance still significantly lags behind that of the similarly sized general-purpose Qwen2.5-7B-Instruct, which achieved a score of 74.35. This leads us to question whether the doctor ranking task demands a higher level of medical knowledge or general logical reasoning abilities. Fine-tuning or continued pre-training of LLMs on medical datasets can enhance their adaptability to medical contexts [44], such as medical consultations. However, the effect on their overall generalizability remains unclear. Further research is necessary to fully harness the medical knowledge embedded within medical LLMs and enhance their specialized reasoning abilities for healthcare-related ranking tasks.
Figure 6: Distribution of marginal probability $p_{k}$ of each graded relevance label in the aforementioned 5-level setting for query-doctor pairs with different ground-truth labels on the DrRank dataset. “NR” is an abbreviation for “Not Relevant”.
5.2.2 Scoring and Explaining Efficiency
To analyze the efficiency of our framework, we present the scoring time and explanation time for an average batch, calculated after removing the two highest and two lowest values, across different models in Table 1.  [^8] Please note that the scoring time of Qwen2.5-32B-Instruct may have been impacted by unknown hardware factors, resulting in possible statistical anomalies. Besides, MING-MOE-7B utilizes the PEFT library for LoRA-based inference, which may introduce additional time delays. We observed that the scoring time of LLMs is significantly slower than that of much smaller cross-encoder models, particularly when the parameter scale of the LLMs exceeds 7 billion. In Section 4.2.2, we introduce the forced elicitation mechanism. This mechanism not only facilitates the elicitation of the desired graded labels but also prevents the one-by-one next-token prediction generation paradigm from affecting the scoring results produced by LLMs, making their scoring time comparable to the prefill time. This also explains why the ranking criteria do not negatively impact the scoring time of the models. However, the computational overhead resulting from an increase in model parameter scale is essentially unavoidable. Additionally, the explanation time of LLMs is inherently longer, and incorporating the ranking criteria further extends this duration. Nevertheless, considering that recommendation explanations can be generated asynchronously for display in real-world scenarios, a longer generation time is generally acceptable.
5.3 Discussions
5.3.1 Intuitive Score Distribution
We group the query-doctor pairs in the DrRank dataset by different ground-truth relevance labels. For every group, we denote $p_{k}$ as the random variable on our graded label $l_{k}$ of the marginal probability $prob_{i,k}$ derived for different query-doctor pairs $(q,{d_{i}})$. We then plot the estimated distribution of $p_{k}$ for each graded label $l_{k}$ respectively. In Fig. 6, we plotted the score distribution of our framework implemented by the Qwen2.5-7B-Instruct model. We observed that the distributions of $p_{k}$ of graded label “Not Relevant”, “Low”, and “Mid” shift down towards 0 as the ground truth relevance increases. Meanwhile, the distributions of $p_{k}$ across relevance label “Top” shift up towards 1. This intuitively reveals how our LLM-based pointwise ranker with fine-grained labels works in practice.
Table 2: The performance of using varying numbers of fine-grained relevance labels on the DrRank dataset. “NR” is an abbreviation for “Not Relevant”. The best and second-best performances are highlighted in bold.
Fine-grained Relevance Labels Criteria NDCG@10 PNR
“Top”, “High”, “Mid”, “Low”, “NR” w/ 74.35 5.31
“Top”, “High”, “Mid”, “Low”, “NR” w/o 72.68 4.9
“High”, “Mid”, “Low”, “NR” w/ 73.57 5.19
“High”, “Mid”, “Low”, “NR” w/o 70.73 4.29
“High”, “Low”, “NR” w/ 66.4 4.58
“High”, “Low”, “NR” w/o 56.35 3.78
“High”, “NR” w/ 63.03 4.71
“High”, “NR” w/o 57.71 6.36
5.3.2 Number of Relevance Labels
We investigate the effect of utilizing varying numbers of fine-grained labels on the performance of the Qwen2.5-7B-Instruct model. As shown in Table 2, increasing the number of relevance labels results in improved performance. This enhancement could be attributed to the fine granularity division in the DrRank dataset, which necessitates providing the model with a maximum number of fine-grained labels to facilitate differentiation. Additionally, we observed that the ranking criteria consistently demonstrate positive effects across different numbers of fine-grained labels. It is noteworthy that in certain scenarios where the NDCG@10 scores are relatively low, the PNR pairwise metric may surprisingly yield high values. For instance, in the 2-level label setting, we observed a PNR value of 6.36, whereas the NDCG@10 score was merely 57.71. This underscores the fundamental difference between the overall optimal nature of the PNR metric and the local top-k optimal focus of the NDCG@10 metric.
Table 3: The performance of each model using different strategies for deriving ranking scores on the DrRank dataset. In each experimental group of strategies, the best performance is highlighted in bold.
Model Strategy Criteria NDCG@10 PNR
Qwen2.5-7B-Instruct $SUM$ w/ 74.35 5.31
Qwen2.5-7B-Instruct $MAX_{logit}$ w/ 75.58 5.64
Qwen2.5-7B-Instruct $MAX_{prob}$ w/ 74.13 5.42
Qwen2.5-14B-Instruct $SUM$ w/ 75.63 6.87
Qwen2.5-14B-Instruct $MAX_{logit}$ w/ 74.82 5.21
Qwen2.5-14B-Instruct $MAX_{prob}$ w/ 75.7 6.9
Qwen2.5-32B-Instruct $SUM$ w/ 77.41 6.13
Qwen2.5-32B-Instruct $MAX_{logit}$ w/ 78.56 5.93
Qwen2.5-32B-Instruct $MAX_{prob}$ w/ 77.38 6.19
5.3.3 Ranking Score Derivation
In our proposed framework, we assign scores to graded labels and calculate the expected value to derive the final score, which we name $SUM$. As described in [11], another straightforward strategy involves using solely the logit of the highest relevance label, termed “Top”, which we denote as $MAX_{logit}$. Alternatively, we can employ the marginal probability of the highest relevance label after applying the softmax function as (2), which we refer to as $MAX_{prob}$. Table 3 presents the performance of various models when using different strategies for deriving ranking scores. We found that the models’ performance fluctuates slightly across three scoring strategies, yet remains consistently high. This suggests that the improvement from scoring fine-grained relevance labels cannot be simply attributed to the expected summation approach, which is essentially equivalent to multiple sampling. This conclusion aligns with the findings reported in [11]. Future work should focus on developing more effective scoring strategies to further enhance the ranking performance of LLMs.
5.3.4 Fairness Consideration
The doctor recommendation system represents a highly sensitive application domain, which can help patients make important and life-changing decisions. It is crucial to ensure that these decisions do not exhibit discriminatory behavior towards specific groups or populations [45]. To assess the fairness of our proposed framework, we evaluate it from three perspectives: diseases, patient gender, and patient regions. For disease fairness, we selected 10 diseases with adequate numbers of candidate doctors. Through proportional sampling based on ground truth labels, we obtained 25 doctors for each disease, ensuring 5 doctors per ground truth label. This approach helps to standardize the difficulty of doctor ranking across diseases as much as possible. We computed the standard deviation of the NDCG@10 metric among the various diseases to measure the model’s bias. This process was repeated 1,000 times, and the average of these standard deviations was taken as the final result. To evaluate fairness related to patient gender and regions, we prepended sensitive information such as “I am male/female” or “I come from an urban/rural area” to the query. If the model displays discriminatory tendencies based on gender or region, these biases will be reflected in the NDCG@10 metric. Table 4 demonstrates the analysis results for fairness in diseases, patient gender, and patient regions. We observed that our LLM-based framework exhibits better disease fairness compared to the bge-reranker-v2-m3 model. Furthermore, the ranking criteria enhance fairness in various disease scenarios by providing standardized guidelines. Meanwhile, we found that our framework and the bge-reranker-v2-m3 model perform similarly in terms of fairness regarding patient gender and regions. However, both of them consistently exhibit stronger biases in terms of gender. Further research is needed on ensuring the fairness of LLMs in ranking tasks.
Table 4: The analysis results for fairness in diseases, patient gender, and patient regions on the DrRank dataset. “SD” is an abbreviation for “standard derivation”. The LLMs with ranking criteria are marked with *.
Model Disease Gender Region
SD NDCG@10 NDCG@10
male female urban rural
bge-reranker-v2-m3 0.0773 67.22 65.5 61.93 62.18
Qwen2.5-7B-Instruct 0.0689 72.31 73.02 73.73 74.39
Qwen2.5-7B-Instruct* 0.0672 74.14 72.56 74.95 75.43
5.3.5 Interpretability Consideration
Our framework improves the interpretability of doctor recommendation by incorporating ranking criteria as standard evaluation guidelines and by generating evaluation rationales to serve as recommendation explanations. Although LLMs excel in text generation and logical reasoning, they occasionally produce hallucinations [46], which undermine the rationality and interpretability of medical advice. To assess our ranking criteria and evaluation rationales, we specifically engaged three medical annotators. The ranking criteria were evaluated based on fluency and professionalism. “Fluency” refers to the LLM-generated text being smooth, coherent, and without grammatical or logical errors. “Professionalism” indicates that the LLM-generated ranking criteria comprehensively reflect doctors’ professional perspectives on a specific disease-treatment pair. The evaluation rationales were assessed based on fluency (as above) and rationality. “Rationality” signifies that the LLM-generated evaluation rationales are medically logical and reasonable. All assessments were scored on a 0 to 5 point scale, with higher scores indicating higher recognition. We used the Qwen2.5-7B-Instruct model to generate these criteria and rationales. In total, 38 ranking criteria for various disease-treatment pairs, and 140 evaluation rationales for 28 disease-treatment pairs (with 5 doctors each) were annotated. Table 5 presents the average scores for each medical annotator, showing that the scores for ranking criteria and evaluation rationales were consistently high across all evaluation perspectives. This clearly demonstrates the great potential of our framework in effectively addressing the interpretability challenge for doctor recommendation in real-world scenarios.
Table 5: The annotation results for interpretability in ranking criteria and evaluation rationales. “Prof.” is an abbreviation for “professionalism”. “a1”, “a1”, and “a3” are refer to the three medical annotators.
Fluency Prof. / Rationality
A1 A2 A3 A1 A2 A3
Ranking Criteria 4.85 3.95 4.97 4.56 4.69 3.87
Evaluation Rationales 4.63 4.65 4.44 3.87 4.39 4.23
6 Conclusion
In this work, we propose a zero-shot LLM-based pointwise ranking framework for doctor recommendation. We conducted a series of comprehensive experiments to validate the accuracy, fairness, and interpretability of our framework. However, our framework relies on LLMs for doctor recommendation, which restricts its applicability in resource-constrained environments. Moreover, the quality of the generated ranking criteria is highly dependent on the quality of the prompts, which requires careful crafting. Future research could aim to address these limitations by exploring more efficient model compression techniques [47] and automated prompt optimization methods [48] for LLM-based doctor recommendation.